OpenStudy (callisto):

Complex number #1 Find the set of complex number such that |\(\frac{z-3}{z+3}\)|=2 How to start?

OpenStudy (anonymous):

If I remember correctly then: \[\Large z\cdot z^*=|z| \] Might want to complex conjugate anyway, for the denominator's sake.

OpenStudy (callisto):

\[z^*\]is the conjugate?

OpenStudy (callisto):

If so, then it should be \[z\cdot z^* = |z|^2\]

OpenStudy (anonymous):

Yes you're right @Callisto, I forgot the square.

OpenStudy (callisto):

z = x+yi \[|z -3| = r_1=\sqrt{(x-3)^2+y^2}\]\[|z +3| = r_2=\sqrt{(x+3)^2+y^2}\] \[|z-3| = |r_1e^{i\theta_1}|\]\[|z+3| = |r_2e^{i\theta_2}|\] \[|\frac{z-3}{z+3}|=\frac{r_1}{r_2}\] Then, some substitution work, we get (x+5)^2 + y^2 = 16 So, the set is {z : z=x+yi, (x+5)^2 + y^2 = 16} Right? Wrong? Any other methods?

OpenStudy (inkyvoyd):

http://www.wolframalpha.com/input/?i= |%28z-3%29%2F%28z%2B3%29|%3D2 looks like it's right